# Are Fact-Checking Tools Helpful? An Evaluation of Google Fact Check

### v3

@@ Abstract
Google Fact Check is a promising search engine that facilitates fact-checking to combat misinformation on social media, especially during significant events such as the COVID-19 pandemic and the U.S. presidential elections. However, its usability has not been thoroughly studied. We evaluated its performance by analyzing the search results regarding 1,000 COVID-19-related false claims and found that only 15.8% of them obtained fact-checking results, and the results returned by the search engine are relatively reliable. We also found that the false claims receiving different fact-checking verdicts (i.e., “False”, “Partly False”, “True”, and “Unratable”) tend to reflect diverse emotional tones, and fact-checking sources tend to check the claims in different lengths and using dictionary words to various extents. Claim variations addressing the same issue yet described differently are likely to obtain disparate fact-checking results. Based on these, we suggested that the quantities of fact-checking results could be optimized and that slightly adjusting input claim wording may be the best practice for users to obtain more useful information. This study aims to contribute to understanding fact-checking tools and promote information integrity.
Index Terms:
@@
communication in support of development; fact-checking; misinformation; evaluation; data analysis; information integrity; search engine; application
I Introduction
@@
Fact-checking has been one of the most common ways to combat misinformation on social media in recent years [1]. It is a significant approach to protect information integrity under the growing influence of social media where people can freely share information and potentially promote misinformation [2]. Several mainstream social media platforms such as Facebook [3] and YouTube [4] have started displaying labels disclosing the potential inaccuracy or other contextual information to foster transparency by collaborating with fact-checking organizations that detect suspicious claims, compile evidence, and publish fact-checking reports. Traditional manual scrutiny usually suffers potential delays and the gap between public focus and fact-checking efforts while facing fast and large-scale dissemination of online information [5], regarding which researchers proposed automated fact-checking frameworks collecting existing fact-checking results or related evidence for real-time evaluation [6].
@@
Neither manual nor automatic approaches can guarantee all false claims potentially bringing significant social influence are fact-checked. When encountering seemingly suspicious claims without contextual information, regular users may struggle to seek related information by themselves: to look for any existing fact-checking results of a claim, they may not have a solid idea of which source has already reviewed the claim and thus end up randomly checking several websites (e.g., PolitiFact, Snopes, and FactCheck.org). In that regard, they may use Google Search, which could return excessive information, such as fact-checking reports and other unhelpful information, making it hard to resolve the concern. Another challenge is that if a claim has been fact-checked by different sources, their potentially different fact-checking verdict terminologies (e.g., “False”, “Mostly False”, “Misleading”, and “Mixed”) and criteria could result in confounding conclusions. For instance, a source could rate a false claim as “Misleading”, while another could give a “Mostly True” rating. Although some researchers compared the fact-checking results from a few popular sources and found them less likely to review the same claims [7], it is unclear whether it is a general case for more sources.
@@
A promising tool to address these two challenges is Google Fact Check, a Google-based search engine for fact-checking results [8]. Users can search for related fact-checking results with complete claims or keywords. Each result on the results page contains a claim assessed by a source, a source name, a fact-checking verdict, a publication date, and a link to the original webpage. Fig. 1 shows the result structure with an example. Besides, an API is also available for software development, such as automated fact-checking. Even though such a tool can significantly improve the efficiency of fact-checking, it is important to validate its ability to provide sufficient and reliable results before rolling it out. Besides, Google Fact Check is an ideal platform to compare the potentially varied fact-checking verdicts and criteria adopted by different sources, such as how many claims were fact-checked by multiple sources and whether their fact-checking results are congruent. To the best of our knowledge, even though researchers contributed to addressing the two challenges mentioned above by, for instance, indicating the limited ability of Google Fact Check to handle complex claims [9] and different types of misinformation [10] and analyzing the fact-checking results from two or three sources [11, 12], no studies have thoroughly studied the performance of Google Fact Check or compared fact-checking results from many sources on a general basis. Therefore, in this study, we aim to understand the performance of Google Fact Check regarding the quality of results in the dimensions that have not been fully explored, such as how relevant the results are to the input claims and the potential relationships between input claims and the fact-checking results they obtained. Specifically, we focused on the following questions:
@@
    To what extent are the fact-checking results on Google Fact Check relevant to the input claims?
@@
    To get the best-matched fact-checking results, is there any relationship between the linguistic characteristics of input claims (e.g., length, emotional tone, and analytical thinking level) and fact-checking results (e.g., the number of results and the fact-checking verdicts)?
@@
    If two or more claims address the same issue but are described differently, to what extent are their fact-checking results congruent?
@@
To conduct this study, we randomly selected 1,000 false claims about COVID-19 from the FakeCovid, a dataset consisting of over 5,000 fact-checking results related to COVID-19 [13], and obtained the corresponding fact-checking results via Google Fact Check API. Then, we analyzed the returned data, including descriptive statistics (i.e., the number of results for each claim, the number of fact-checking verdicts, and source reliability) and inferential statistics regarding the correlation between the linguistic characteristics of the input claims and their fact-checking results. We found that 842 (84.2%) of the 1,000 false claims did not obtain any fact-checking results from Google Fact Check, and the remaining 158 (15.8%) received at least one result each. They received 290 results in total, among which 94.46% were relevant to the input claims, i.e., the claims assessed by fact-checkers and the input claims focused on the same topic. Among these relevant results, 91.54% were rated “False” or “Partly False” by the sources with high reliability. As for the relationships between claims and fact-checking results, the original claims receiving different verdicts (i.e., “False”, “Partly False”, “True”, and “Unratable”) tend to reflect diverse emotional tones, and the claims reviewed by different sources tend to be in various lengths and use dictionary words to various extents. We also found that claim variations addressing the same issue are likely to obtain distinctive fact-checking results. We suggested that the quantities of fact-checking results could be optimized and that slightly adjusting input claim wording may be the best practice for users to obtain more useful information. This study aims to contribute to the understanding of fact-checking tools and promote information integrity.
@@
The rest of this paper is organized as follows. In Section II, we introduce our methodology, including the data collection process and data analysis methods. Then, we present and discuss the results in Section III. Finally, we summarize our findings, limitations, and future work in Section IV.
II Methodology
II-A Data Collection
@@
To study the performance of Google Fact Check, we randomly selected 1,000 false claims about COVID-19, a hot topic that aroused numerous rumors and conspiracies during the pandemic, from the FakeCovid [13], a dataset consisting of over 5,000 COVID-19-related fact-checking results from Poynter and Snopes, two reputable fact-checking sources, where the former is also the leader of the International Fact-Checking Network (IFCN), the biggest fact-checking alliance joined by more than 100 fact-checking organizations globally [14].
@@
We programmed via Google Fact Check API to obtain the fact-checking results for these claims. The returned metadata were originally in JSON fields, including text, claimant, publisher, textual rating, and review date [15]. We renamed the fields to make them more readable: Input Claim, Claim Assessed by Fact-Checkers (i.e., the real claim fact-checked, which could be identical to or different from the input claim), Fact-Checking Source, Fact-Checking Verdict, Publication Date, and URL (see Fig. 1 for an example). All the data we analyzed in this study were obtained in December 2022 via the API except the input claims from the FakeCovid dataset. The results are shown in Section III-A.
II-B Data Sanitization
II-B1 Relevance of Fact-Checking Results to Input Claims
@@
Since Google Fact Check is essentially a search engine, a claim assessed by fact-checkers could overlap with the input claim to different extents: (1) they are identical, i.e., the exact claim has been fact-checked; (2) the input claim has yet to be reviewed, but some other relatable claims are returned, so the relevance of the returned item is worth an examination. For instance, when the claim “Queen Elizabeth II is infected with the new coronavirus” has not been fact-checked yet, the fact-checking results for another claim “Queen Elizabeth died because of the COVID-19 vaccine” could be returned, even though the latter deviates from the point “infection”. To investigate to what extent the returned results are relevant to their original input claims, we recruited three coders from our research team to individually rate whether each claim assessed by fact-checkers is relevant to the corresponding input claim, i.e., whether they address the same issue. As a result, 262 (90.66%) of the 289 fact-checking results were rated unanimously by three coders (i.e., “relevant” or “irrelevant”), indicating a high agreement among coders, even though the Krippendorff’s alpha is K = 0.476 [16]. Based on that, we rated a result “relevant” to the input claim if at least two of the three coders agreed. The rating results are shown in Section III-A.
II-B2 Mapping of Fact-Checking Verdict Terminology
@@
Each fact-checking result on Google Fact Check contains a verdict indicating the accuracy of a claim. Since Google Fact Check is a search engine collecting fact-checking results from a wide range of sources adopting different verdict terminologies (e.g., “Mostly False”, “Mixture”, and “Mostly Inaccurate”) and rating criteria, synonymous verdicts (e.g., “Mostly False” and “Partly True”) could bring a challenge to our analysis. We reviewed the verdict definitions adopted by all the sources involved in the returned results and summarized them as follows: if there is clear evidence, the verdict is essentially “false”, “true”, or a mixture of both; otherwise, the accuracy is inconclusive. Thus, we mapped the original verdicts into four categories: “False”, “Partly False”, “True”, and “Unratable”. Specifically, we directly mapped four verdicts (i.e., “False”, “Partly False”, “True”, and “Unratable”) to the categories under the same name and the other verdicts to these four categories based on the definitions adopted by their sources. [^1] The complete datasets and documentation are available on OSF: https://osf.io/zkbd4/?view_only=804bdd91c7f340eeacc3bf3494c06930
@@
For the verdicts in a long sentence instead of simple words, the three coders in Section II-B1 manually reviewed and mapped them to the proper categories the same way they coded in Section II-B1. The resulting Krippendorff’s alpha is K = 0.817, indicating a high inter-coder agreement [16]. For any disagreements among the coders, we accepted the choice by at least two coders; if they all chose different categories, they had an open discussion until reaching an agreement. As a result, all the original verdicts were mapped to these four categories for better analysis. Although it is impossible to cover all the fact-checking sources in the world and the verdict terminologies they adopted, these four categories can theoretically cover any verdict from any source. The analysis results of the fact-checking verdicts are shown in Section III-B.
II-C Source Reliability
@@
II-D Relationship Between Claims and Results
@@
II-D1 Number of Results vs. Characteristics of Input Claims
We tested whether the linguistic characteristics of input claims are likely to result in obtaining different numbers of fact-checking results. Since the variables do not follow normal distributions, we calculated Spearman’s rank correlation coefficients ranging from -1 to 1, where the positive and negative signs indicate the directions of correlation and a greater deviation from 0 indicates a stronger correlation [20]. The results are shown in Section III-D1.
II-D2 Relevance of Results vs. Characteristics of Input Claims
In Section II-B1, three coders reviewed the fact-checking results and evaluated whether each claim assessed by fact-checkers was relevant to the original input claim, i.e., whether they focused on the same topic. We wondered whether such relevance is related to the characteristics of input claims. To do so, we grouped the fact-checking results we obtained based on whether they were rated relevant or irrelevant in Section II-B1 and performed the Kruskal-Wallis H test for any significant difference in the characteristics of input claims. The results are shown in Section III-D2. We performed the Kruskal-Wallis H test because the “relevance” is nominal data and should not be calculated as meaningful numbers. It is the same for the calculations in the following Sections II-D3 and II-D4.
II-D3 Fact-Checking Verdicts vs. Characteristics of Input Claims
In Section II-B2, the fact-checking verdicts from various sources were mapped to four categories (i.e., “False”, “Partly False”, “True”, and “Unratable”) for better analysis. We wondered whether claims with different characteristics are likely to obtain varied verdicts, e.g., claims showing a lower critical thinking level are more likely to be rated “False”. To do so, we first excluded the irrelevant and non-English fact-checking results coded in Section II-B1. Then, we grouped the remaining results based on the four categories and performed the Kruskal-Wallis H test to detect any significant difference in the characteristics of input claims. The results are shown in Section III-D3.
II-D4 Fact-Checking Sources vs. Characteristics of Input Claims
In Section II-C, we evaluated the reliability of the fact-checking sources. Some sources may have a taste for specific topics or claims in specific linguistic patterns, which could drive them to adopt various fact-checking methodologies and criteria. To better understand the potential relationship between sources and claim characteristics, we performed the Kruskal-Wallis H test to detect the significant difference in the characteristics of claims across the sources. The results are presented in Section III-D4.
II-E Input Claim Variation

### v5

# Are Fact-Checking Tools Helpful? An Exploration of the Usability of Google Fact Check

@@ Abstract
Fact-checking-specific search tools such as Google Fact Check are a promising way to combat misinformation on social media, especially during events bringing significant social influence, such as the COVID-19 pandemic and the U.S. presidential elections. However, the usability of such an approach has not been thoroughly studied. We evaluated the performance of Google Fact Check by analyzing the retrieved fact-checking results regarding 1,000 COVID-19-related false claims and found it able to retrieve the fact-checking results for 15.8% of the input claims, and the rendered results are relatively reliable. We also found that the false claims receiving different fact-checking verdicts (i.e., “False,” “Partly False,” “True,” and “Unratable”) tend to reflect diverse emotional tones, and fact-checking sources tend to check the claims in different lengths and using dictionary words to various extents. Claim variations addressing the same issue yet described differently are likely to retrieve distinct fact-checking results. We suggest that the quantities of the retrieved fact-checking results could be optimized and that slightly adjusting input wording may be the best practice for users to retrieve more useful information. This study aims to contribute to the understanding of state-of-the-art fact-checking tools and information integrity.
Keywords:
@@
Fact-checking Misinformation Infodemic Search engine Data analysis User experience Information integrity.
1 Introduction
@@
Fact-checking has been one of the most common ways to combat misinformation on social media in recent years [23]. It is a significant approach to protect information integrity under the growing influence of social media where people can freely share information and potentially disseminate misinformation [9]. Several mainstream social media platforms such as Facebook [1] and YouTube [2] display labels disclosing potential inaccuracy or context to foster transparency by collaborating with fact-checking organizations that regularly detect suspicious claims, compile evidence, and publish fact-checking reports. Traditional manual scrutiny usually suffers potential delays and the gap between public focus and fact-checking efforts while facing fast and large-scale dissemination of online information [20], regarding which researchers proposed automated fact-checking frameworks to collect existing fact-checking results or related evidence for real-time review [12].
@@
Neither manual nor automatic approaches can guarantee all false claims potentially bringing great social impact are fact-checked. When encountering seemingly suspicious claims without contextual information, regular users may struggle to seek related information by themselves: to look for any existing fact-checking results of a claim, they may not have a solid idea of which source has already reviewed the claim and end up randomly checking several websites (e.g., PolitiFact, Snopes, and FactCheck.org). In that regard, they may use Google Search, which could return excessive information, such as fact-checking reports and other unhelpful information, making it hard to resolve the concern. Another challenge is that if a claim has been fact-checked by multiple sources, their potentially different fact-checking verdict terminologies (e.g., “False,” “Mostly False,” “Misleading,” and “Mixed”) and rating criteria could result in confounding conclusions. For instance, regarding an identical false claim, a verdict could be rated “misleading” by one source while “mostly true” by another. Although some researchers compared the fact-checking results from a few popular sources and found them less likely to review the same claims [16], whether it is a general case for most sources is unclear.
@@
A promising tool to address these two challenges (i.e., lack of useful tools to integrate fact-checking results and insufficient knowledge of the congruence of the fact-checking results rendered by various sources) is Google Fact Check Tools, a Google-based search engine for fact-checking results [3]. Users can search for fact-checking results by complete claims or keywords. Each result on the results page contains a claim assessed by a source, a source name, a fact-checking verdict, a publication date, and a URL to the original report. Fig. 1 shows the result structure in an example. An API is also available for software development, such as automated fact-checking. Although such a tool can significantly improve the efficiency of fact-checking, it is important to validate its ability to provide sufficient and reliable results before rolling it out. Besides, Google Fact Check is an ideal platform to compare the potentially varied fact-checking verdicts across sources, such as how many claims have been fact-checked by multiple sources and whether their fact-checking results are congruent. To the best of our knowledge, Google Fact Check is the only search engine specifically for fact-checking as of this study. Even though researchers contributed to addressing the aforementioned challenges by, for instance, indicating the limited ability of Google Fact Check to handle complex claims [18] and different types of misinformation [14] and analyzing the fact-checking results rendered by a few sources [17, 19], no studies have thoroughly explored the performance of Google Fact Check or similar fact-checking-specific search engines, nor did they compare fact-checking results across sources on a general basis. Therefore, in this study, we aim to understand the usability of fact-checking-specific search tools by evaluating the performance of Google Fact Check regarding the quality of results in the dimensions that have not been fully explored, such as how relevant the retrieved fact-checking results are to the input claims and their potential correlations, so that we may shed light on the best practices for users to retrieve useful information. Specifically, we focused on the following questions:
@@
    To what extent are the fact-checking results retrieved by Google Fact Check relevant to the input claims?
@@
    Is there any correlation between the linguistic characteristics of input claims (e.g., length, emotional tone, analytical thinking level) and the retrieval of the best-matched fact-checking results?
@@
    If multiple claims address the same issue but are described differently, to what extent are their fact-checking results congruent?
@@
In general, we made the following contributions:
  1. 1.
    We evaluated the performance of Google Fact Check by analyzing the retrieved fact-checking results regarding 1,000 COVID-19-related false claims. In our experiments, 842 (84.2%) of the claims did not retrieve any fact-checking results, and the remaining 158 (15.8%) retrieved at least one result each. In total, 290 results were returned, among which 94.46% were relevant to the input claims, i.e., the claims scrutinized by fact-checkers and the input claims addressed the same topic. Among these relevant results, 91.54% were rated “false” or “partly false” by fact-checking sources with high reliability.
  2. 2.
    We explored the correlations between claims and fact-checking results via data analyses. The original claims receiving different verdicts (i.e., “False,” “Partly False,” “True,” and “Unratable”) tend to reflect diverse emotional tones, and the claims scrutinized by different sources tend to be in various lengths and use dictionary words to various extents.
  3. 3.
    We found that claim variations addressing the same issue are likely to obtain distinctive fact-checking results, shedding light on the best practices for users to retrieve the most useful information.
@@
The rest of this paper is organized as follows. In Section 2, we introduce our methodology, including the data collection process and data analysis methods. Then, we present our results in Section 3. Finally, we discuss our findings, limitations, and future work in Section 4.
2 Methodology
2.1 Data Collection
@@
Misinformation can be found anywhere, and fact-checking can be applied to any topic. In this study, we decided to focus on COVID-19 as it was one of the representative topics arousing numerous rumors and conspiracies during the pandemic, even resulting in “infodemic,” [10], and there are sufficient datasets available for our experiments. We leveraged the FakeCovid [21], a dataset consisting of over 5,000 COVID-19-related fact-checking results from Poynter and Snopes, two reputable fact-checking sources, where the former is also the leader of the International Fact-Checking Network (IFCN), the biggest fact-checking alliance joined by more than 100 fact-checking organizations globally [4]. This dataset was compiled in the early stage of the pandemic when knowledge about COVID-19 was lacking, and related misinformation was rampant on social media. This multilingual dataset covers multiple domains of COVID-19, such as origin, spread, treatment, and conspiracy [21]. To study the performance of Google Fact Check, we excluded non-English results from the FakeCovid dataset and randomly selected 1,000 false claims for our experiments.
@@
We programmed via Google Fact Check API to obtain the fact-checking results for these claims. The returned metadata were originally in JSON fields, including text, claimant, publisher, textual rating, and review date [7]. We renamed the fields to make them more readable: Input Claim, Claim Assessed by Fact-Checkers (i.e., the real claim fact-checked, which could be identical to or different from the input claim), Fact-Checking Source, Fact-Checking Verdict, Publication Date, and URL (see Fig. 1 for an example). All the data we analyzed in this study were obtained in December 2022 via the API except the input claims from the FakeCovid dataset. The results are shown in Section 3.1.
2.2 Data Sanitization
2.2.1 Relevance of Fact-Checking Results to Input Claims
@@
Since Google Fact Check is essentially a search engine, a claim assessed by fact-checkers could overlap with the input claim to different extents: (1) they are identical, i.e., the exact claim has been fact-checked; (2) the input claim has yet to be reviewed, but some other relatable claims are returned, so the relevance of the returned item is worth an examination. For instance, when the claim “Queen Elizabeth II is infected with the new coronavirus” has not been fact-checked yet, the fact-checking results for another claim “Queen Elizabeth died because of the COVID-19 vaccine” could be returned, even though the latter deviates from the point “infection.” To investigate to what extent the returned results are relevant to their original input claims, we recruited three coders from our research team to individually rate whether each claim assessed by fact-checkers is relevant to the corresponding input claim, i.e., whether they address the same issue. As a result, 262 (90.66%) of the 289 fact-checking results were rated unanimously by three coders (i.e., “relevant” or “irrelevant”), indicating a high agreement among coders, even though the Krippendorff’s alpha is K = 0.476 [15]. Based on that, we rated a result “relevant” to the input claim if at least two of the three coders agreed. The rating results are shown in Section 3.1.
2.2.2 Mapping of Fact-Checking Verdict Terminology
@@
Each fact-checking result on Google Fact Check contains a verdict indicating the accuracy of a claim. Since Google Fact Check is a search engine collecting fact-checking results from a wide range of sources adopting different verdict terminologies (e.g., “Mostly False,” “Mixture,” and “Mostly Inaccurate”) and rating criteria, synonymous verdicts (e.g., “Mostly False” and “Partly True”) could bring a challenge to our analysis. We reviewed the verdict definitions adopted by all the sources involved in the returned results and summarized them as follows: if there is clear evidence, the verdict is essentially “false,” “true,” or a mixture of both; otherwise, the accuracy is inconclusive. Thus, we mapped the original verdicts into four categories: “False,” “Partly False,” “True,” and “Unratable.” Specifically, we directly mapped four verdicts (i.e., “False,” “Partly False,” “True,” and “Unratable”) to the categories under the same name and the other verdicts to these four categories based on the definitions adopted by their sources. [^1] The complete datasets and documentation are available on OSF: https://osf.io/zkbd4/?view_only=804bdd91c7f340eeacc3bf3494c06930
@@
For the verdicts in a long sentence instead of simple words, the three coders in Section 2.2.1 manually reviewed and mapped them to the proper categories the same way they coded in Section 2.2.1. The resulting Krippendorff’s alpha is K = 0.817, indicating a high inter-coder agreement [15]. For any disagreements among the coders, we accepted the choice by at least two coders; if they all chose different categories, they had an open discussion until reaching an agreement. As a result, all the original verdicts were mapped to these four categories for better analysis. Although it is impossible to cover all the fact-checking sources in the world and the verdict terminologies they adopted, these four categories can theoretically cover any verdict from any source. The analysis results of the fact-checking verdicts are shown in Section 3.2.
2.3 Source Reliability
@@
2.4 Correlation Between Claims and Results
@@
2.4.1 Number of Results vs. Characteristics of Input Claims
We tested whether the linguistic characteristics of input claims are likely to result in obtaining different numbers of fact-checking results. Since the variables do not follow normal distributions, we calculated Spearman’s rank correlation coefficients ranging from -1 to 1, where the positive and negative signs indicate the directions of correlation and a greater deviation from 0 indicates a stronger correlation [22]. The results are shown in Section 3.4.1.
2.4.2 Relevance of Results vs. Characteristics of Input Claims
In Section 2.2.1, three coders reviewed the fact-checking results and evaluated whether each claim assessed by fact-checkers was relevant to the original input claim, i.e., whether they focused on the same topic. We wondered whether such relevance is related to the characteristics of input claims. To do so, we grouped the fact-checking results we obtained based on whether they were rated relevant or irrelevant in Section 2.2.1 and performed the Kruskal-Wallis H test for any significant difference in the characteristics of input claims. The results are shown in Section 3.4.2. We performed the Kruskal-Wallis H test because the “relevance” is nominal data and should not be calculated as meaningful numbers. It is the same for the calculations in the following Sections 2.4.3 and 2.4.4.
2.4.3 Fact-Checking Verdicts vs. Characteristics of Input Claims
In Section 2.2.2, the fact-checking verdicts from various sources were mapped to four categories (i.e., “False,” “Partly False,” “True,” and “Unratable”) for better analysis. We wondered whether claims with different characteristics are likely to obtain varied verdicts, e.g., claims showing a lower critical thinking level are more likely to be rated “False.” To do so, we first excluded the irrelevant and non-English fact-checking results coded in Section 2.2.1. Then, we grouped the remaining results based on the four categories and performed the Kruskal-Wallis H test to detect any significant difference in the characteristics of input claims. The results are shown in Section 3.4.3.
2.4.4 Fact-Checking Sources vs. Characteristics of Input Claims
In Section 2.3, we evaluated the reliability of the fact-checking sources. Some sources may have a taste for specific topics or claims in specific linguistic patterns, which could drive them to adopt various fact-checking methodologies and criteria. To better understand the potential relationship between sources and claim characteristics, we performed the Kruskal-Wallis H test to detect the significant difference in the characteristics of claims across the sources. The results are presented in Section 3.4.4.
2.5 Input Claim Variation